# Matthias von Davier

### v1

@@ 3 Kish’s (1965) Effective Sample Size Estimate
Consider a random variable $Y$ with finite mean $\mu$ and finite variance $\sigma^{2}$. Let $y_{i}\in\mathbb{R},i=1,\dots,N$ denote the observed values in a sample of size $K$ and assume the sample requires weighting with weights $w_{k}$. Assume we are studying a weighted estimator of the sample mean using weights $w_{k}\in\mathbb{R}^{+}$, that is
@@
Another scenario is that the $y_{i}$ are statistics based on smaller samples taken from certain well-defined groups (sampling units, sub-populations, etc.). In that case, $y_{k}=f_{k}\left(y_{k1},\dots,y_{kN_{k}}\right)$ are (potentially also weighted) aggregates based on the within group observations $y_{k1},\dots,y_{kN_{k}}$.
@@
    The quality of the variance estimator $Var_{w}\left(y\right)$, may differ from the quality of the unweighted estimate $Var_{1}\left(y\right).$ The term quality is chosen to be deliberately vague here.
@@
$Var\left(\overline{y}_{w}\right)=Var\left(\frac{\sum_{k=1}^{K}w_{k}y_{k}}{\sum_{k=1}^{K}w_{k}}\right)=\frac{\sum_{k=1}^{K}w_{k}^{2}Var\left(y_{k}\right)}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}$
@@
$Var\left(\overline{y}_{w}\right)=\frac{\sum_{k=1}^{K}w_{k}^{2}\sigma^{2}}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}=\sigma^{2}\frac{\sum_{k=1}^{K}w_{k}^{2}}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}=\frac{\sigma^{2}}{N}\frac{\overline{w^{2}}}{\overline{w}^{2}}$
@@ 5.2 Total Variance under Multiple Imputations
5.3 Welch Test
The modified Welch (1947) test uses the d.f. $\nu_{1}$ and $\nu_{2}$ of each of the variances of two samples to pool the variance. The Welch-Satterthwaite formula used in this case can be adjusted using the improved approximation proposed here. Then we have

### v2

@@ 3 Kish’s (1965) Effective Sample Size Estimate
Consider a random variable $Y$ with finite mean $\mu$ and finite variance $\sigma^{2}$. Let $y_{i}\in\mathbb{R},i=1,\dots,N$ denote the observed values in a sample of size $K$ and assume the sample requires weighting using the weights $w_{k}$. We are studying a weighted estimator of the sample mean using weights $w_{k}\in\mathbb{R}^{+}$, that is
@@
Another scenario is that the $y_{i}$ are statistics based on smaller samples taken from certain well-defined groups (sampling units, sub-populations, etc.). In that case, $y_{k}=f_{k}\left(y_{k1},\dots,y_{kN_{k}}\right)$ are (potentially also weighted) aggregates based on the within group observations $y_{k1},\dots,y_{kN_{k}}$. In this case, each of the groups may receive a different weight to account for and correct discrepancies between expected population proportion vs. sample sizes for each group.
@@
    The quality of the variance estimator $Var_{w}\left(y\right)$, may differ from the quality of the unweighted estimate $Var_{1}\left(y\right).$ The term quality is chosen to be deliberately vague here. One well known issue with weighting is that it improve one quality (reduce bias) at the cost of another (increase uncertainty).
@@
$Var\left(\overline{y}_{w}\right)=Var\left(\frac{\sum_{k=1}^{K}w_{k}y_{k}}{\sum_{k=1}^{K}w_{k}}\right)=\frac{\sum_{k=1}^{K}w_{k}^{2}Var\left(y_{k}\right)}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}},$
@@
$Var\left(\overline{y}_{w}\right)=\frac{\sum_{k=1}^{K}w_{k}^{2}\sigma^{2}}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}=\sigma^{2}\frac{\sum_{k=1}^{K}w_{k}^{2}}{\left(\sum_{k=1}^{K}w_{k}\right)^{2}}=\frac{\sigma^{2}}{N}\frac{\overline{w^{2}}}{\overline{w}^{2}}.$
@@ 5.2 Total Variance under Multiple Imputations
5.3 A Corrrected Welch Test $d.f.$
The modified Welch (1947) test uses the d.f. $\nu_{1}=N_{1}-1$ and $\nu_{2}=N_{2}-1$ of each of the variances of two samples to pool the variance. The Welch-Satterthwaite formula used in this case can be adjusted using the corrected Satterthwaite formula derived above. Then we have